maria has $3.00 in nickles, dimes and quarters. If she has twice as many dimes as quarters and five more nickels than dimes, how many coins of each does she have?
Answer by stanbon(75887)

Jakathy is 32 years older than her nephew, Eric. Five years ago, her age was seventeen times as old as his. What is Jakathy's age?
------------
Equations:
j = e + 32
(j-5) = 17(e-5)
-----------------------
Substitute for "j" and solve for "e":
(e+32-5) = 17e-85
e+27 = 17e-85
16e = 112
e = 7 yrs old (erik's age now)
-----
Since j = e+32, j=39 yrs (Jakathy's age now)
============================
Cheers,
Stan H.